OpenStudy (anonymous):

I'm having trouble with an inverse function problem in Cal 2: Find the value of the derivative of the inverse of the function. Given that f(x)=4^(sqrt(x))-x+3, find f-1'(3). I can't figure out what to do with the constant raised to the power of sqrt(x)

OpenStudy (anonymous):

The easiest way to find an inverse of a function is to simply swap the x and the y (or in this case "f(x)") and then solve for f(x). Is it this: \[4^{\sqrt{x}}-x-3=f(x)?\] or this: \[4^{\sqrt{x}-x-3}=f(x)?\]

OpenStudy (anonymous):

oops I meant +3 not -

OpenStudy (anonymous):

Boy that's nastier than it looks.